OpenStudy (anonymous):

How would you find the inverse of (e^x)/(1+e^x)?

OpenStudy (anonymous):

Solve for x: y = e^x/(1 + e^x) (1 + e^x) * y = e^x ye^x + y = e^x y = e^x - ye^x y = e^x( 1 - y) y/(1-y) = e^x ln(y/(1-y)) = x swap the variables around to make it look more consistent: g(x) = ln(x/(1-x)) f(x) = e^x/(1+e^x) g(x) is the inverse of f(x)